40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 12 17:37:11 2011, Jenne, Update, SUS, ETMY hopefully good again 
    Reply  Fri Aug 12 20:28:13 2011, Jenne, Update, SUS, ETMY hopefully good again 
       Reply  Sat Aug 13 02:31:42 2011, kiwamu, Update, SUS, Re: ETMY hopefully good again 
          Reply  Mon Aug 15 13:16:50 2011, Jenne, Update, SUS, Re: ETMY hopefully good again 
Message ID: 5237     Entry time: Mon Aug 15 13:16:50 2011     In reply to: 5221
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Re: ETMY hopefully good again 

Quote:

I guess the ETMY suspension is still fine. Their OSEM DC voltage and the free swinging spectra look healthy.

It could be a failure in the initial guess for fitting.

Quote from #5216

I'm no longer convinced that ETMY is healthy. I can't fit the peaks to get the input matrix.

 Turns out I was missing a critical step in the process...running makeSUSspectra.m  After I do that, everything is back under control, and ETMY looks fine. 

I'm almost done doing the peak-fitting and matrix inversion for all optics.

ELOG V3.1.3-